# FAIR Universe HiggsML Uncertainty Challenge Competition

### v3

@@ Abstract
The FAIR Universe – HiggsML Uncertainty Challenge focuses on measuring the physics properties of elementary particles with imperfect simulators due to differences in modelling systematic errors. Additionally, the challenge is leveraging a large-compute-scale AI platform for sharing datasets, training models, and hosting machine learning competitions. Our challenge brings together the physics and machine learning communities to advance our understanding and methodologies in handling systematic (epistemic) uncertainties within AI techniques.
fair-universe@lbl.gov
https://fair-universe.lbl.gov
Compiled: September 7, 2025
@@ 1.1 Background and impact
For several decades, the discovery space in almost all branches of science has been accelerated dramatically due to increased data collection brought on by the development of larger, faster instruments. More recently, progress has been further accelerated by the emergence of powerful AI approaches, including deep learning, to exploit these data. However, an unsolved challenge that remains, and must be tackled for future discovery, is how to effectively quantify and reduce uncertainties, including understanding and controlling systematic uncertainties (also named epistemic uncertainties in other fields). This is widely true across scientific and industrial applications involving measurement instruments (medicine, biology, climate science, chemistry, and physics, to name a few). A compelling example is found in analyses to further our fundamental understanding of the universe through analysis of the vast volumes of particle physics data produced at CERN, in the Large Hadron Collider (LHC).
@@
Ten years ago, part of our team co-organised the Higgs Boson Machine Learning Challenge (HiggsML) pmlr-v42-cowa14 , the most popular Kaggle challenge at the time attracting 1785 teams. This challenge has significantly heightened interest in applying Machine Learning (ML) techniques within High-Energy Physics and, conversely, has exposed physics issues to the ML community. Whereas previously, the most effective methods predominantly relied on boosted decision trees, Deep Learning has since gained prominence (see, e.g., HEP ML living review). While the LHC has not (yet) discovered new physics beyond the Higgs boson, it has accumulated vast data and will continue to accumulate more data well into the next decade. There is a discovery potential in very precise measurements of particle properties, in particular the Higgs boson which can only be studied at the LHC. This demands meticulous measurement of uncertainties and their incorporation into developing sophisticated AI techniques.
@@
High energy physics (HEP) relies on statistical analysis of aggregated observations. Therefore, the interest in uncertainty-aware ML methods in HEP is nearly as old as the application of ML in the field. Advanced efforts that integrate uncertainties into the ML training include approaches that explicitly depend on nuisance parameters Cranmer:2015bka ; Baldi:2016fzo ; Brehmer:2019xox ; Brehmer:2018hga ; Brehmer:2018kdj ; Brehmer:2018eca ; Nachman:2019dol ; Ghosh:2021roe , that are insensitive to nuisance parameters Blance:2019ibf ; Englert:2018cfo ; Louppe:2016ylz ; Dolen:2016kst ; Moult:2017okx ; Stevens:2013dya ; Shimmin:2017mfk ; Bradshaw:2019ipy ; ATL-PHYS-PUB-2018-014 ; DiscoFever ; Wunsch:2019qbo ; Rogozhnikov:2014zea ; 10.1088/2632-2153/ab9023 ; clavijo2020adversarial ; Kasieczka:2020pil ; Kitouni:2020xgb ; Estrade:2019gzk ; Ghosh:2021hrh , that use downstream test statistics in the initial training Wunsch:2020iuh ; Elwood:2020pik ; Xia:2018kgd ; deCastro:2018mgh ; Charnock_2018 ; Alsing:2019dvb ; Simpson:2022suz ; Feichtinger:2021uff ; Layer:2023lwi , and that use Bayesian neural networks for estimating uncertainties Kasieczka:2020vlh ; Bollweg:2019skg ; Araz:2021wqm ; Bellagente:2021yyh . Many of these topics were covered in a forward-looking review-type article in Ref. Chen:2022pzc . However, these developments all report technique performance on different ad-hoc datasets, so it is difficult to compare their merits. This competition aims at providing a common ground, with a dataset of sufficient complexity, equipped with systematic bias parameterisations, and a metric.
We aim to address the issue of systematic uncertainties within a specific domain. Yet, the techniques developed by the challenge participants will apply to identifying, quantifying, and correcting systematic uncertainties in other areas. A critical aspect of this work is to provide a dataset and platform that conforms to the principles of findability, accessibility, interoperability, and reusability - inspiring the name of our project: FAIR Universe. We plan to keep our submission platform accessible even after the challenge concludes, establishing it as a lasting benchmark. This initiative should significantly influence research in uncertainty-aware AI/ML techniques, which currently suffer from a critical shortage of datasets and benchmarks dedicated to their research and development.
@@ 1.2 Novelty
This entirely new public competition builds on our experience running several competitions in particle physics and broader. These include the original HiggsML data challenge pmlr-v42-cowa14 , the TrackML Challenges (NeurIPS 2018 competition) TrackMLAccuracy2019 ; TrackMLThroughput , the LHC Olympics lhc-olympics , AutoML/AutoDL automl ; autodl , and other competitions. Building on the foundation of the original HiggsML challenge, this competition introduces a significant change by using simulated data that includes biases (or systematic effects) in the test dataset. In addition, participants are asked to provide a confidence interval and not just a point estimate. We have developed an innovative metric to assess their performance.
While there have been previous challenges focusing on meta-learning and transfer-learning, such as the NeurIPS 2021 and 2022 meta-learning challenges elbaz2021pmlr ; carriónojeda2022neurips22 , Unsupervised and Transfer Learning guyon2012analysis , challenges related to bias e.g. Crowd bias challenge crowd-bias-challenge , and those addressing distribution shifts, like the Shifts challenge series, and CCAI@UNICT 2023 ccaiunict-2023 , to the best of our knowledge, this is the first challenge that requires participants to handle systematic uncertainty.
Moreover, this project is connecting the Perlmutter system at NERSC, a large-scale supercomputing resource featuring over 7000 NVIDIA A100 GPUs, with Codabench, a new version of the renowned open-source benchmark platform Codalab competitions. Our challenge is among the first to utilise this platform for a public competition, marking a significant milestone in accessible, high-performance computing for AI research.
@@ 2 Data
We are using a simulated particle physics dataset for this competition to produce data representative of high energy proton collision data collected by the ATLAS experiment 2008atlas at the Large Hadron Collider (LHC) Evans_2008 . The dataset is created using two widely used simulation tools, Pythia 8.2 Sjostrand:2014zea and Delphes 3.5.0 deFavereau:2013fsa . We have organised the dataset into a tabular format where each row corresponds to a collision event with 28 features, the measurements recorded from a single proton bunch crossing of interest (see Appendix A). The detailed list of features is available in Appendix C, preceded by a brief introduction to special relativity that can be used to understand them better Appendix B. The events are divided into two categories (see Table 1): signal and background. The signal category includes collision events with a Higgs boson decaying into tau pairs (see Figure 1) (one decaying into a lepton, the other one into a hadronic tau), while the background category includes other processes (subcategories) leading to a similar final state.
Figure 1: Diagram of the particles in the final state chosen: one lepton, one tau hadron, up to two jets, and the missing transverse momentum vector, see Appendix A for details.
Due to its complexity, the process of generating events is computationally intensive; use of the Perlmutter supercomputer allowed us to create a vast amount of data, about 280 million events, which is almost three orders of magnitude larger than for the HiggsML competition. It will be made publicly available under the Creative Commons Attribution license to serve as a benchmark dataset after the competition.
@@
In addition, we have developed a biasing script capable of manipulating a dataset by introducing six parameterised distortions as a function of six Nuisance Parameters [^1] The name Nuisance Parameter, commonly used in the physics literature, refers to a parameter governing a specific parameterisation of a systematic bias. Nuisance Parameters can be in part constrained from the data itself but the name implies that constraining them is only interesting as an auxiliary task in the process of determining a parameter of interest like the signal strength $\mu$. (the systematic biases); see details in Appendix D. For example, a detector miscalibration can cause a bias in other features in a cascade way, or in another case, the magnitude of a particular background (e.g. the $t\bar{t}$) contribution can change so that the composition of the background (thus the feature distributions) can be different. In both cases, the inference would be done on a dataset not i.i.d to the training dataset.
To avoid any leakage, the dataset, with all labels, was handled on the private disks of just two team members before being uploaded to Codabench, where it is secured. The random generator seeds, which could be used to reproduce the dataset, are also secured.
Table 1: Summary of the dataset for each category and subcategory. "Number Generated" is the number of events available. In contrast, "LHC events" is the average number in this category in a pseudo-experiment corresponding to running of the Large Hadron Collider for 10 fb-1, corresponding to approximately 800 billion inelastic proton collisions, or 2 weeks in summer 2024 conditions.
Process Number Generated LHC Events Label
Higgs 52101127 1015 signal
Z Boson 221724480 1002395 background
Di-Boson 2105415 3783 background
$t\bar{t}$ 12073068 44190 background
@@ 3 Tasks and application scenarios
The participant’s objective is to develop an estimator for the number of Higgs boson events in a dataset analogous to results from Large Hadron Collider experiments. Such a measurement is typical to those carried over at the Large Hadron Collider, which allows us to strengthen (or invalidate!) our understanding of the fundamental laws of nature.
@@
The primary metric is the signal strength ($\mu$), which is the number of estimated Higgs boson events divided by the number of such events predicted by the Standard Model, which is the reference theory. The challenge involves estimating $\mu$’s true value, $\mu_{true}$, which may vary from one (in practice in the range 0.1 to 3) and is inherently unknown.
@@
Participants are tasked with generating a 68.27% Confidence Interval (CI) for $\mu$, incorporating both aleatoric (random) and epistemic (systematic) uncertainties rather than a single-point estimate. The implementation of the six different systematic uncertainties is detailed in Appendix D.
@@
The primary simulation dataset assumes a $\mu$ of one. Participants receive a training subset, where events are labelled based on their event type (e.g. Higgs boson event). The organisers provide the script to generate unlabelled pseudo-experiment datasets from the primary simulation dataset for any value of $\mu$ and the six systematic biases. The participant’s model should be able to reverse the process and provide a 68.27% CI on $\mu$ for any pseudo-experiment.
@@
In a machine learning context, the task resembles a transduction problem with distribution shift: it requires constructing a $\mu$ interval estimator from labelled training data and biased unlabelled test data. One possibility is to train a classifier to distinguish Higgs boson from the background, with robustness against bias achieved possibly through data augmentation (or an adversarial approach, or black box optimization or any other novel approach) via the provided script.
@@
This challenge shifts focus from the qualitative discovery of individual Higgs boson events (which was the focus of our first challenge) to the quantitative estimation of overall Higgs boson counts in test sets, akin to assessing disease impact on populations rather than diagnosing individual cases.
4 Metrics
Participants must submit a model to the Codabench platform that can analyse a dataset to determine $(\mu_{16},\mu_{84})$, which represents the bounds of the 68.27% Confidence Interval (CI) for $\mu$.
The model’s performance is assessed based on two criteria:
  • Precision: The narrowness of the CI (narrower is preferable).
  • Coverage: The accuracy of the CI in reflecting the measurement’s uncertainty, meaning there should be a 68.27% probability that $\mu_{\rm truth}$ falls within the CI.
The model that predicts $(\mu_{16},\mu_{84})$ serves as an estimator. In statistics, the effectiveness of an estimator is not evaluated based on a single case but rather over numerous instances. Therefore, the model undergoes evaluation across a series of $N_{\text{test}}$ pseudo-experiments, each characterized by randomized biases and $\mu_{\text{truth}}$, to assess it against the specified criteria. We introduce a novel uncertainty metric, which we call Coverage Score. This metric is divided into two components:
The first component is the Average Interval Width $w$, calculated as follows:
$w=\frac{1}{N_{test}}\sum_{i=1}^{N}|\mu_{84,i}-\mu_{16,i}|\;,$ (1)
where $\mu_{16,i}$ and $\mu_{84,i}$ are the bounds of the 68.27% CI for each experiment $i$ within the range $1,N_{\text{test}}]$.
The second component $c$ quantifies the frequency with which the true value of $\mu_{\rm true}$ falls within the 68.27% Confidence Interval (CI), as illustrated in Figure 2:
$c=\frac{1}{N_{test}}\sum_{i=1}^{N}\ 1\ \textrm{if}\ \mu_{\textrm{true},i}\in[\mu_{16,i}-\mu_{84,i}]\;.$ (2)
@@ 4 Metrics
Figure 2: Coverage plot: all the predicted intervals (blue lines) for each pseudo experiment generated for a given $\mu_{\rm true}$ (vertical dotted line). The coverage (here $70\pm 5\%$) is determined by the fraction of time the horizontal blue lines intersect the vertical line. The average width of the interval is here 1.068.
If the confidence interval $\mu_{16,i},\mu_{84,i}]$ accurately represents the 68.27% quantile, the true value of $\mu$ should lie within this interval in 68.27% of the pseudo-experiments. Consequently, we employ a penalising function $f$ that penalises models that deviate from this 68.27% reference (Figure 3) :
$c\in[0.6827-2\sigma_{68},0.6827+2\sigma_{68}]$ $:f(c)=1$ (3)
$c<0.6827-2\sigma_{68}$ $:f(c)=1+\left|\frac{c-(0.6827-2\sigma_{68})}{\sigma_{68}}\right|^{4}$ (4)
$c>0.6827+2\sigma_{68}$ $:f(c)=1+\left|\frac{c-(0.6827+2\sigma_{68})}{\sigma_{68}}\right|^{3}$ (5)
Figure 3: Penalising function as a function of the coverage value $c$. The shape of the function is determined by uncertainty on the coverage evaluation, governed by $N_{\textrm{t}est}$, the number of pseudo-experiments. A larger $N_{\textrm{t}est}$ leads to a narrower function.
Here, $\sigma_{68}$ is the standard deviation for a binomial distribution with $p=0.6827$, given by
$\sigma_{68}=\sqrt{\frac{(1-0.6827)0.6827}{N_{\textrm{test}}}}~,$ (6)
@@
this parameter establishes a tolerance zone around the anticipated value of $0.6827$, accommodating statistical variances inherent in the evaluation process. We opted for an asymmetric penalty function because, within the field of High Energy Physics (HEP), overestimating uncertainty is deemed more acceptable than underestimating it Nachman:2012zf ; Nachman:2014hsa . Hence, coverage exceeding 68.27% incurs a lesser penalty than coverage falling below 68.27%.
The final Coverage Score used to rank participants is calculated as follows:
@@
$w$ represents the average width of the Confidence Interval, $c$ is the coverage parameter, and $\epsilon=10^{-2}$ is a regularization term to guard against submissions that report unrealistically narrow CIs. Using the $\ln()$ function ensures that higher scores are awarded to superior submissions and that the score variations remain within a modest range.
@@
To ensure efficient use of resources and timely updates to the leaderboard, each participant’s model inference is executed across 100 pseudo-experiments times five distinct values of $\mu_{\rm truth}$. In the Final phase of the competition, the best submission from each participant will be evaluated over 1,000 pseudo-experiments, times ten different $\mu_{\rm truth}$ values, to minimize random noise in the evaluation.
5 Baselines, code, and material provided
A Starting Kit is available on the challenge website. This kit includes code for installing necessary packages, loading and visualising data, training and evaluating a model and preparing a submission for the competition. This should facilitate local testing and adjustments by participants before their official submissions and allow them to train and evaluate models. Moreover, the competition adheres to a specific interface, details of which are documented in the Starting Kit and demonstrated through the provided sample code.

### v4

# FAIR Universe HiggsML Uncertainty Dataset and Competition

@@ Abstract
The FAIR Universe – HiggsML Uncertainty Challenge focused on measuring the physical properties of elementary particles with imperfect simulators. Participants were required to compute and report confidence intervals for a parameter of interest regarding the Higgs boson while accounting for various systematic (epistemic) uncertainties. The dataset is a tabular dataset of 28 features and 280 million instances. Each instance represents a simulated proton-proton collision as observed at CERN’s Large Hadron Collider in Geneva, Switzerland. The features of these simulations were chosen to capture key characteristics of different types of particles. These include primary attributes, such as the energy and three-dimensional momentum of the particles, as well as derived attributes, which are calculated from the primary ones using domain-specific knowledge. Additionally, a label feature designates each instance’s type of proton-proton collision, distinguishing the Higgs boson events of interest from three background sources. As outlined in this paper, the permanent release of the dataset allows long-term benchmarking of new techniques. The leading submissions, including Contrastive Normalising Flows and Density Ratios estimation through classification, are described. Our challenge has brought together the physics and machine learning communities to advance our understanding and methodologies in handling systematic uncertainties within AI techniques.
@@ 1.1 Background and impact
For several decades, the discovery space in almost all branches of science has been accelerated dramatically due to increased data collection brought on by the development of larger, faster instruments. More recently, progress has been further accelerated by the emergence of powerful AI approaches, including deep learning, to exploit this data. However, an unsolved challenge that remains, and must be tackled for future discovery, is how to effectively quantify and reduce uncertainties, including understanding and controlling systematic uncertainties (also named epistemic uncertainties in other fields). A compelling example is found in analyses to further our fundamental understanding of the universe through analysis of the vast volumes of particle physics data produced at CERN, in the Large Hadron Collider (LHC) [1]. Ten years ago, part of our team co-organised the Higgs Boson Machine Learning Challenge (HiggsML [2, 3], the most popular Kaggle challenge at the time attracting 1785 teams. This challenge has significantly heightened interest in applying Machine Learning (ML) techniques within High-Energy Physics (HEP) and, conversely, has exposed physics issues to the ML community. Whereas previously, the most effective methods predominantly relied on boosted decision trees, Deep Learning has since gained prominence (see, e.g., HEP ML living review [4]). While the LHC has not (yet) discovered new physics beyond the Higgs boson, it has accumulated vast data and will continue to accumulate more data well into the next decade. There is a discovery potential in very precise measurements of particle properties, particularly of the Higgs boson.
@@
High-energy physics relies on statistical analysis of aggregated observations. Therefore, the interest in uncertainty-aware ML methods in HEP is nearly as old as the application of ML in the field. Advanced efforts that integrate uncertainties into the ML training include approaches that explicitly depend on nuisance parameters [5, 6, 7, 8, 9, 10, 11, 12, 13, 14], that are insensitive to nuisance parameters [15, 16, 17, 18, 19, 20, 21, 22, 23, 24, 25, 26, 27, 28, 29, 30, 31, 32], that use downstream test statistics in the initial training [33, 34, 35, 36, 37, 38, 39, 40, 41, 42, 43], and that use Bayesian neural networks for estimating uncertainties [44, 45, 46, 47]. Many of these topics were covered in recent forward-looking review-type articles in Refs. [48, 49]. However, these developments all report technique performance on different ad-hoc datasets, so it is difficult to compare their merits. The Fair Universe HiggsML Uncertainty Challenge, an official NeurIPS 2024 competition, aimed to provide a common ground, with a dataset of sufficient complexity, equipped with systematic bias parameterisations, and a metric.
We aim to address the issue of systematic uncertainties within a specific domain. Yet, the techniques developed by the challenge participants will apply to identifying, quantifying, and correcting systematic uncertainties in other areas, particularly other science disciplines.
@@ 1.2 Novelty
This entirely new public competition has built on our experience running several competitions in particle physics and beyond. These include the original HiggsML challenge [2], the TrackML Challenges (NeurIPS 2018 competition) [50, 51], the LHC Olympics [52], AutoML/AutoDL [53, 54], and other competitions. Building on the foundation of the HiggsML challenge, this competition introduces a significant change by using simulated data that includes biases (or systematic effects). In addition, participants were asked to provide a confidence interval and not just a point estimate.
While there have been previous challenges focusing on meta-learning and transfer-learning, such as the NeurIPS 2021 and 2022 meta-learning challenges [55, 56], Unsupervised and Transfer Learning [57], challenges related to bias e.g. Crowd bias challenge [58], and those addressing distribution shifts, like the Shifts challenge[59] series, and CCAI@UNICT 2023 [60], this is the first challenge and dataset that requires participants to handle systematic uncertainty. Moreover, this project is connecting the Perlmutter system at NERSC [61], a large-scale supercomputing resource featuring over 7000 NVIDIA A100 GPUs, with Codabench [62], a new version of the renowned open-source benchmark platform Codalab [63, 64]. Due to its complexity, the process of generating events was computationally intensive; use of the Perlmutter supercomputer allowed us to create a vast amount of data – hundreds of millions of events compared to less than a million events for the HiggsML competition, which will serve as a long-lasting benchmark.
@@ 2 Data
The dataset is publicly available on the Zenodo platform [65]. The data is saved as a tabular parquet [66] file of 16 GB and is accompanied by a Croissant JSON metadata file. The dataset comprises 280M simulated proton-proton collision events and is weighted to represent two weeks of LHC data taking. A separate 120M i.i.d dataset has been used for the final results in section 5 and is kept private for future over-training checks.
@@
We are using a simulated particle physics dataset for this competition to produce data representative of high-energy proton collision data collected by the ATLAS experiment [67] at the LHC. The dataset [65] was created with two widely-used simulation tools, Pythia 8.2 [68] and Delphes 3.5.0 [69]; all the configuration and data pre-selection code is available from [70]. This required 1.8 million CPU core hours. We have organised the dataset into a tabular format where each row corresponds to a collision event and each of the 28 columns corresponds to a feature. The detailed dataset description is in Appendix A, Appendix B and Appendix C; it is mostly taken from the public unpublished Fair Universe whitepaper [71] which served as detailed documentation for the competition. Part of the features are primary features, essentially the energy and direction of a small set of particles, and the remainder are derived features, computed from the primary ones with domain knowledge. The events are divided into two categories (see Table 1): signal and background. The signal category includes collision events with a Higgs boson decaying into pairs of tau particles (see Figure 7) (one decaying, into, in addition to neutrino(s), a light lepton, the other one into a set of hadrons hence the name hadronic tau), while the background category includes other processes (subcategories) leading to a similar final state, but without an intermediate Higgs boson.
@@
In addition, we provide a biasing script capable of manipulating a dataset by introducing six parameterised distortions as a function of six corresponding Nuisance Parameters [^1] The name Nuisance Parameter, commonly used in the physics literature, refers to a parameter governing a specific parameterisation of a systematic bias. Nuisance Parameters can be in part constrained from the data itself. Still, the name implies that constraining them is only interesting as an auxiliary task in the process of determining a parameter of interest like the signal strength $\mu$. (the systematic biases); see details in Appendix D. For example, a detector miscalibration can cause a bias in other features in a cascading way, or in another case, the magnitude of a particular background (e.g. the $t\bar{t}$) contribution can change so that the feature distributions can be different. In both cases, the inference would be done on a dataset not i.i.d. to the training dataset.
@@ 3 Tasks and application scenarios
The participant’s objective is to develop an estimator for the number of Higgs boson events in a dataset analogous to results from LHC experiments. Such a measurement is typical of those carried out at the LHC, which allows us to strengthen (or invalidate) our understanding of the fundamental laws of nature.
@@
The primary metric is the signal strength ($\mu$), which is the number of estimated Higgs boson events divided by the number of such events predicted by the Standard Model, which is the reference theory. The challenge involved estimating $\mu$’s true value, $\mu_{true}$, which may vary from one (in practice for the challenge in the range 0.1 to 3) and is inherently unknown.
@@
Participants were tasked with generating a 68.27% Confidence Interval (CI) for $\mu$, incorporating both aleatoric (random) and epistemic (systematic) uncertainties rather than a single-point estimate. The six different systematic uncertainties are implemented in Appendix D.
@@
The primary simulation dataset assumes a $\mu$ of one. Participants receive a training subset, where events are labelled based on their event type (e.g. Higgs boson event). We provide a script to generate unlabelled pseudo-experiment datasets from the primary simulation dataset for any value of $\mu$ and the six systematic biases. The participant’s model should be able to reverse the process and provide a 68.27% CI on $\mu$ for any pseudo-experiment.
@@
In a machine learning context, the task resembles a transduction problem with distribution shift: it requires constructing a $\mu$ interval estimator from labelled training data and biased unlabelled test data. One possibility is to train a classifier to distinguish Higgs boson from the background, with robustness against bias achieved possibly through data augmentation (or an adversarial approach, or black box optimisation or any other novel approach) via the provided script.
@@
This challenge shifts focus from the qualitative discovery of individual Higgs boson events (which was the focus of our first challenge [2]) to the quantitative estimation of overall Higgs boson counts in test sets, akin to assessing disease impact on populations rather than diagnosing individual cases.
3.1 Metrics
Participants provided a model that can analyse a pseudo-experiment to determine $(\mu_{16},\mu_{84})$, the bounds of the 68.27% (approximately one standard deviation of a standard normal distribution) Confidence Interval (CI) for $\mu$. The model is evaluated from the set of $\mu_{16,i},\mu_{84,i}]$ intervals obtained from $N_{\text{test}}$ pseudo-experiments, see 1(a). The model’s performance is assessed based on two criteria: Average Interval Width $w$ (the smaller the better) computed as $w=\frac{1}{N_{test}}\sum_{i=1}^{N}|\mu_{84,i}-\mu_{16,i}|$ and the Coverage, the frequency with which $\mu_{\rm truth}$ is covered by the CI (the closer to the standard 68.27% probability the better) computed as $c=\frac{1}{N_{test}}\sum_{i=1}^{N}\ 1\ \textrm{if}\ \mu_{\textrm{true},i}\in[\mu_{16,i},\mu_{84,i}]$.
A penalising function $f$ has been defined to penalise the departure of $c$ from the expected 68.27%, taking into account $\sigma_{68}=\sqrt{\frac{(1-0.6827)0.6827}{N_{\textrm{test}}}}$ the binomial statistical error on $c$:
$f(c)=1+\mathbb{I}_{c<0.6827-2\sigma_{68}}\cdot\left|\frac{c-(0.6827-2\sigma_{68})}{\sigma_{68}}\right|^{4}+\mathbb{I}_{c>0.6827+2\sigma_{68}}\cdot\left|\frac{c-(0.6827+2\sigma_{68})}{\sigma_{68}}\right|^{3}\,.$ (1)
@@ 3.1 Metrics
Figure 1: (1(a)) Coverage plot: all the predicted intervals (blue lines) for each pseudo experiment generated for a given $\mu_{\rm true}$ (vertical dotted line). The coverage (here $70\pm 5\%$) is determined by the fraction of time the horizontal blue lines intersects the vertical line. (1(b)) Penalising function as a function of the coverage value $c$, for two values of $N_{\textrm{t}est}$, the number of pseudo-experiments.
@@
We opted for an asymmetric penalty function because, within the High Energy Physics (HEP) field, overestimating uncertainty is deemed more acceptable than underestimating it [72, 73]. Hence, coverage exceeding 68.27% incurs a lesser penalty than coverage falling below 68.27%. The final Quantile Score (the larger the better) used to rank participants is calculated as follows:
@@
where $w$ represents the average width of the Confidence Interval, $c$ is the coverage, and $\epsilon=10^{-2}$ is a regularisation term to guard against submissions that report unrealistically narrow CIs. To ensure efficient use of resources, each participant’s model inference was executed across 100 pseudo-experiments times 10 trials, each with distinct values of $\mu_{\rm truth}$, with a time limit of 20s per inference on CPU or GPU. In the Final phase of the competition, each participant’s best submission was evaluated over 100 pseudo-experiments, times 1000 trials, to minimize the statistical variance.
3.2 Limitations
The main limitation of the setup is that biases can be exactly parameterised: we are in the "known unknowns" regime. "Unknown unknowns", unexpected biases, are not covered.
The dataset has been produced using well-known standard software for event generation and detector simulation. However, a proper physics measurement would require more complex software, several orders of magnitude slower, yielding marginally different simulated data. The methods developed on our dataset would perform equally well, provided they are fully retrained.
For each instance of the datasets, the features provided are essentially the energy and direction of a small set of particles, and derived quantities. A real physics measurement may also rely on additional quantities related to the quality of particle identification or to other particles in the same proton-proton collision. Nevertheless, the algorithms developed on our dataset should require limited added complexity to deal with additional features.
4 Software
Alongside the dataset, a GitHub repository [74] with the relevant code for reading and analysing it is made available. This includes a Jupyter notebook starting kit, simple baseline models, a small sample of the dataset, and code to compute the score.